Session: Feb 2016 (session 1)
Wavelength: 5 cm
To give feedback on this experiment, click
(L. Nicotra at Tue Mar 01 06:54:24 2016) Success:
Hartebeesthoek (Jonathan Quick at Tue Mar 01 12:48:08 2016) Success:
Phase reference source and target both set whilst still being observed. No other known problems.
(Ruben Fernández at Tue Mar 01 14:48:48 2016) Success:
No known problem.
(A.Maccaferri at Tue Mar 01 15:40:03 2016) Success:
(Geert Kuper at Tue Mar 01 16:52:18 2016) Success:
(Uwe at Thu Mar 03 13:07:29 2016) Minor failures: HIGH
The LCP problem is back. Tsys in LCP as high as 200K from 7:30 to 11:30 UT. What is new is that the problem occurred and vanished within an observation. The receiver will be inspected in the lap after the session.
(Carlo Migoni at Fri Mar 04 10:00:14 2016) Success:
Onsala (Jun Yang at Sat Mar 05 11:24:17 2016) Success:
No known problems.
(Pawel Wolak at Sat Mar 05 13:13:25 2016) Success:
(Vladislavs Bezrukovs at Wed Mar 23 08:56:15 2016) Success: